OpenStudy (anonymous):

What is the sum of the arithmetic sequence 140, 133, 126…, if there are 33 terms?

OpenStudy (anonymous):

\[Sn=(a1+an)*n/2\] an=a1+(n-1)*r r=7

OpenStudy (anonymous):

d here is -7. Therefore, we get: a33 = 140 + 32(-7) = -84. So, we substitute in the Sn formula and we get: Sn = (140 - 84)*33/2 = 924. If I did not make any careless mistake.

OpenStudy (anonymous):

thankss